show that the points A= (1,5) B= (2,4) C= (-3,5) lie on a circle with center (-1,2). What is the radius of this circle?
Please and thank you

Answer by stanbon(75887)

show that the points A= (1,5) B= (2,4) C= (-3,5) lie on a circle with center
(-1,2). What is the radius of this circle?
----
Show that (-1,2) is the same distance for A, from B, and from C
------------
(-1,2) to A::: sqrt((1--1)^2+(2-5)^2) = sqrt[4+9] = sqrt(13)
-----
(-1,2) to B::: sqrt[(-1-2)^2+(2-4)^2] = sqrt(9 + 4) = sqrt(13)
-----
(-1,2) to C::: sqrt[(-3--1)^2 + (5-2)^2] = sqrt(16 + 9) = sqrt(25) = 5
-----
They are not on a circle with center at (-1,2)
-----
Cheers,
Stan H.
=============